OpenStudy (anonymous):

compute the following linear combinations (79.4,58.1)+3(-0.8,6.3)

OpenStudy (saifoo.khan):

let vector a = (79.4,58.1) = 79.4+58.1j and vector b = 3(-0.8+6.3j) = 2.4+18.9j (vector a) + (vector b) = 81.8+77j

OpenStudy (anonymous):

so the answer is (81.8, 77j) ??

OpenStudy (saifoo.khan):

right.